Penn State Fleet which operates and manages car rentals for Penn State employees found that the tire lifetime for their vehicles has a mean of 50,000 miles and standard deviation of 3500 miles.

What is the probability that the sample mean lifetime for these 50 vehicles exceeds 51,000?
Answer #1

Solution :

Given that,

mean = \mu = 50000

standard deviation = \sigma =3500

n=50

\mu\bar x = \mu =50000

\sigma\bar x = \sigma / \sqrt n = 3500/ \sqrt50 = 494.9748

P(\bar x >51000 ) = 1 - P(\bar x < 51000)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (51000-50000) /494.9748 ]

= 1 - P(z < 2.02)

Using z table

= 1 - 0.9783

=0.0217

probability= 0.0217